[Hallc_running] [New Logentry] RC Update

latif at jlab.org latif at jlab.org
Sat Nov 10 13:30:03 EST 2018

Logentry Text:
- The beam was not that great since yesterday swing shift. Yesterday around 4:00 PM, the accelerator had high voltage issue in zone 1L23. This resulted about 3 hours of downtime and beam came back around 7:00 PM. Over the owl shift, the beam tripped frequently. Today the exact same problem happened again around noon. We are expecting another 3-4 hours of downtime.

- We finished all other kinematic settings from group-1. Currently, we are at 1-12, the last setting from group-1.

- We will move to kinematics group-5 once the last setting from group-1 is done.

- At this moment we are still waiting for the beam.

There will be no RC meeting today and tomorrow. I  will be sending out the update.


This is a plain text email for clients that cannot display HTML.  The full logentry can be found online at https://logbooks.jlab.org/entry/3624028
-------------- next part --------------
An HTML attachment was scrubbed...
URL: <https://mailman.jlab.org/pipermail/hallc_running/attachments/20181110/cbfdcfb2/attachment.html>

More information about the Hallc_running mailing list